OpenStudy (anonymous):

@AlexandervonHumboldt2 help me lol

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

\[\frac{ 3 }{ 8 }\div \frac{ 2 }{ 3 }+\frac{ 1 }{ 4 }\] @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

okm so 3/8 / 2/3=3/8 * 3/2

OpenStudy (alexandervonhumboldt2):

(3*3)/(8*2)=???

sammixboo (sammixboo):

Yes, that equals that due to the CCFM method

OpenStudy (anonymous):

9/16 @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

now 9/16+1/4= can you guess anything?

OpenStudy (alexandervonhumboldt2):

make their demnominator equal: 16=4*4 so it will be 9/16+4/16=??????

OpenStudy (anonymous):

13/16

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

okay then then

OpenStudy (anonymous):

okay then whats else @AlexandervonHumboldt2

OpenStudy (anonymous):

@sammixboo

sammixboo (sammixboo):

That's it!

OpenStudy (anonymous):

so that's the answer :) 13/16

sammixboo (sammixboo):

Right =)